How do you graph this? (x-4)^2+(y+6)=16
Answer by lwsshak3(11628)

How do you graph this? (x-4)^2+(y+6)=16
**
(x-4)^2+(y+6)=16
rewrite:
(x-4)^2+y+6=16
y= -(x-4)^2+10
This is an equation of a parabola of the standard form:y=-A(x-h)^2+k, with (h,k) being the (x,y) coordinates of the center. The negative coefficient A means the parabola opens downward and the parabola has a maximum:
For given equation:
vertex:(4,10)
x-intercepts
set y=0
-(x-4)^2+10=0
(x-4)^2=10
(x-4)=±√10≈±3.16
x≈4±3.16=7.16 and 0.84
you now have 3 points with which you can draw the graph: (4,10), (7.16,0) and (0.84,0)